OpenStudy (anonymous):

Justify the first step that David took in verifying that x = 5 is a solution for the equation 8x – 13 = 2x + 17. 8x – 13 = 2x + 17 8(5) – 13 = 2(5) + 17 A. David multiplied 8 by 5. B. David subtracted 13 from both sides. C. David replaced x with 5. D. David added 17 to both sides

OpenStudy (sleepyhead314):

well, looking at the steps 8x – 13 = 2x + 17 8(5) – 13 = 2(5) + 17 what does it look like he did?